OpenStudy (anonymous):

Mike packed an equal number of flashlights in each of x gift boxes. The number of compartments in each gift box was 3 more than the number of boxes. The number of flashlights in each compartment of a box was 2 more than the number of compartments in each box. The expression below shows the total number of flashlights that Mike packed in all the boxes: x(x + 3)(x + 5) Which expression shows the number of flashlights in each box? x + 3 x + 5 (x + 3)(x + 5) x(x + 3)(x + 5) my choice is b @swamijatin2 @perl

OpenStudy (swamijatin2):

o find your answer, you have to distribute, there is none other to able to distribute than d. I will find more complex information if this doesn't work.

OpenStudy (anonymous):

N it's not asking for that

OpenStudy (anonymous):

it's asking for the number of flashlights in each box, and it says "The number of compartments in each gift box was 3 more than the number of boxes. The number of flashlights in each compartment of a box was 2 more than the number of compartments in each box."

OpenStudy (perl):

flashlights per box = total number of flashlights / total number of boxes

OpenStudy (anonymous):

Yes .

OpenStudy (perl):

there are a total of x(x+3)(x+5) flashlights. there are a total of x boxes

OpenStudy (perl):

what is x(x+3)(x+5)/x

OpenStudy (anonymous):

OOow it would be c

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@iGreen ???

OpenStudy (anonymous):

@TheSmartOne @kidrah69

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

seriously

OpenStudy (anonymous):

I don't want google, I want live help that's the point of OS is it not?

TheSmartOne (thesmartone):

Yes, C would be correct. c:

OpenStudy (anonymous):

so please don't answer my question if you don't know what I am asking, thanks and thanks tso do you know anyone else who can give an opinin?

TheSmartOne (thesmartone):

well Perl kinda already did. He told you to do: \(\sf\Large\frac{ x(x+3)(x+5)}{x }=(x+3)(x+5)\)

OpenStudy (anonymous):

Ok I was confused about that for a second but now i see it better

TheSmartOne (thesmartone):

(: